OpenStudy (anonymous):

Kevin buys a motorcycle for $8,000. The dealer is charging him an annual interest rate of 9.25% and is using the add-on method to compute monthly payments. If he pays off the loan in 84 months, what are his monthly payments? If he makes a down payment of $2,400, how much will his monthly payments be? Do not round until the final answer. Then, round to the nearest cent. A: $156.90; $109.83 B: $156.90; $220.13 C: $314.48; $109.83 D: $156.90; $204.41

OpenStudy (anonymous):

I get answer A. 8,000(.0925)(7)=5,180 in interest charges So (5,180+8,000)/84= 156.90 (8,000-2,400)(.0925)(7)=3,626 So (8,000-2,400 +3,626)/84= 109.83

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

were did the 7 come from in the first part